A point charge (q = +3.5 micro-C) is moved between two points in an electric field. If this requires 7.0 x 10-5 J of work, what is the potential difference between the two points?


1
Expert's answer
2021-01-13T11:37:24-0500

By definition, the potential difference between the two points is equal to the work required to move a unit charge from one point to another. Since it requires "W = 7\times 10^{-5}J" to move the charge "q = 3.5\times 10^{-6}C", then for moving a unit charge the work will be:


"U = \dfrac{W}{q} = \dfrac{7\times 10^{-5}J}{3.5\times 10^{-6}} = 20V"

Answer. 20V.